# Joint Processing and Transmission Energy Optimization for ISAC in Cell-Free Massive MIMO with URLLC

### v2

@@ V-B GOPS Analysis of Digital Operations for Sensing at the Cloud
In this subsection, we present the GOPS analysis for sensing operations, including signal transmission and processing the received signal. The GOPS corresponding to the sensing transmissions consists of computing the sensing precoding vector $\textbf{w}_{0}$, denoted by $C_{\rm prec-comp}^{\rm s}$, and obtaining the sensing signal by multiplying the sensing symbols by the sensing precoding vector and the sensing power coefficient. Multiplying the sensing symbols by the sensing precoding vector and the sensing power coefficient costs $12L_{d}MN_{\rm tx}$ real multiplications/divisions. The unitary matrix U in (24) (subspace spanned by the UE channel estimation vectors) is already obtained when computing the RZF precoding vectors by matrix inversion and the corresponding $\mathbf{L}\mathbf{D}\mathbf{L}^{H}$ decomposition [36, Lem. B.2]. Therefore, the computational complexity of the ZF precoding vector $\textbf{w}_{0}$ is given by
@@
After transmitting the signal in downlink, the reflected signals along with the interference signals are received at the receiver APs and sent to the cloud. At the cloud, these signals are processed for a specific sensing application. In this work we consider target detection and assess the number of real multiplications/divisions required to compute the test statistics. For target detection problems, we usually compute the test statistics and compare them with a threshold. The target is declared detected if the value of the test statistics is greater than the threshold. We assume that the threshold is constant. Therefore, we can neglect the computational complexity of obtaining the threshold. However, test statistics should be obtained for each transmission. Hence, the sensing GOPS is obtained as
@@
where $C_{\rm se-comp}$ denotes the computational complexity for computing the detector test statistics. Given the clutter-unaware detector in (17), the computational complexity is computed as
@@
where we used the block diagonal structure of C in getting the inverse of it.
The computational complexity for each step of computing the test statistics of clutter-aware detector in (18) are listed in Table I.
TABLE I: Computational complexity with clutter-aware MAPRT detector
Operation Computational complexity
$\textbf{G}[m],\quad\forall m$ $\mathcal{O}\left(20\,L_{d}\,MN_{\rm tx}N_{\rm rx}\right)$
a $\mathcal{O}\left(8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}\right)\,$
b $\mathcal{O}\left(8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}\,\right)$
C $\mathcal{O}\left(4\,L_{d}MN_{\rm rx}\left(N_{\rm tx}^{2}+N_{\rm tx}\right)\right)$
D $\mathcal{O}\left(4L_{d}\,\left((MN_{\rm tx})^{2}+MN_{\rm tx}\right)\right)$
E $\mathcal{O}\left(8\,L_{d}\,M^{2}\,N_{\rm rx}\,N_{\rm tx}^{2}\right)$
$\begin{bmatrix}\textbf{C}&\textbf{E}\\ \textbf{E}^{H}&\textbf{D}\end{bmatrix}^{-1}$ $\mathcal{O}\left(\frac{8}{3}\left(\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{3}-(1+M^{2})N_{\rm tx}N_{\rm rx}\right)\right)$
$\textbf{D}^{-1}$ $\mathcal{O}\left(\frac{8}{3}\left((M^{2}N_{\rm tx}N_{\rm rx})^{3}-M^{2}N_{\rm tx}N_{\rm rx}\right)\right)$
$\begin{bmatrix}\textbf{a}^{H}&\textbf{b}^{H}\end{bmatrix}\begin{bmatrix}(.)&(.)\\ (.)&(.)\end{bmatrix}\begin{bmatrix}\textbf{a}\\ \textbf{b}\end{bmatrix}$ $\mathcal{O}\left(8\,\left(\left((1+M^{2})N_{\rm tx}N_{\rm rx}\right)^{2}+(1+M^{2})N_{\rm tx}N_{\rm rx}\right)\right)$
@@ VI Joint blocklength and Power Optimization
In this paper, we aim to optimize the blocklength and the power control coefficients to minimize the energy consumption while URLLC and sensing requirements are satisfied. In the following subsections, we cast two non-convex optimization problems: i) minimizing only the energy consumption for transmission; ii) minimizing E2E energy consumption by taking into account the energy consumption at the radio site and the cloud. [^2] In addition to considering radio and cloud processing energy consumption, it is important to take fronthaul energy consumption into account when assessing E2E energy consumption. However, in our analysis, where the primary focus is on the interplay between processing and radio resources, we treat fronthaul energy consumption as a fixed component and do not include it in our considerations. The corresponding algorithms to these optimization problems are represented by SeURLLC+ and E2E SeURLLC+, respectively.
VI-A Minimizing Energy Consumption for Transmission
The optimization problem can be cast as
[…]
subject to $\epsilon_{i}\leq\epsilon^{\rm(ub)}_{i}\leq\epsilon^{\rm(th)}_{i},\quad\forall i>0$ (51b)
$L\leq L_{\rm max},\quad$ (51c)
$\overline{\mathsf{SINR}}_{s}\geq\gamma_{\rm s},\quad$ (51d)
$P_{k}\leq P_{\rm tx},\quad\quad k=1,\cdots,N_{\rm tx}$ (51e)
$P_{\rm tr}=\sum_{k=1}^{N_{\rm tx}}P_{k}$ is the average power consumption for transmission, given as $P_{\rm tr}=\sum_{j=0}^{N_{\rm ue}}\rho_{j}=\|\bm{\rho}\|^{2}$, where $\bm{\rho}=[\sqrt{\rho_{0}}\ \ldots\ \sqrt{\rho_{N_{\rm ue}}}]^{T}$, due to the unit-power centralized precoding vectors. Constraints (51b) and (51c) correspond to the URLLC requirements. The $\gamma_{\rm s}$ is the required sensing SINR that is selected according to the target detection performance requirement and $P_{\rm tx}$ is the maximum transmit power per AP.
@@ VI-A Minimizing Energy Consumption for Transmission
$\underset{\bm{\rho},\bm{\chi},\textbf{r}\geq\textbf{0},L>L_{p},\bar{L}>0}{\textrm{minimize}}\quad\frac{\|\bm{\rho}\|^{2}}{\overline{L}}\,$ (52a)
@@
$\underset{\bm{\rho},\bm{\chi},\textbf{r}\geq\textbf{0},L,\bar{L}>L_{p},\chi_{0}\geq 0}{\textrm{minimize}}\quad\frac{\|\bm{\rho}\|^{2}}{\overline{L}}+\lambda\chi_{0}\,,$ (53a)
subject $\textrm{subject to}\quad{L-L_{p}}\leq\frac{2}{\overline{L}^{(c-1)}}-\frac{\overline{L}}{\left(\overline{L}^{(c-1)}\right)^{2}},$ (53b)
@@
The steps of FPP-SCA and CCP procedure are outlined in Algorithm 1. We empirically observed that setting $\overline{L}^{(0)}=\frac{1}{L_{\rm max}-L_{p}}$, $\rho_{0}^{(0)}=0$ and $\sqrt{\rho_{i}}^{(0)}=10^{-3}\sqrt{P_{\rm tx}/N_{\rm ue}}$ for $i>0$ yields satisfactory results.
Algorithm 1 FPP-SCA and CCP Procedure for Problem (52).
1:Initialization: Set arbitrary $\bm{\rho}^{(0)},\textbf{r}^{(0)}\geq\textbf{0}$, the solution accuracy parameters $\epsilon,\epsilon_{\chi}>0$, $\overline{L}^{(0)}=\frac{1}{L_{\rm max}-L_{p}}$, and $\lambda>0$. Set the iteration counter to $c=0$, and a maximum iteration number $c_{\rm max}$. Moreover, let $\mathsf{F}^{(0)}=\infty$ be the initial value of objective function and $\Delta\mathsf{F}^{(c)}=\mathsf{F}^{(c)}-\mathsf{F}^{(c-1)}$ with $\mathsf{F}^{(-1)}=0$.
2:$c\leftarrow c+1$.
3:While $\Delta\mathsf{F}^{(c-1)}\geq\epsilon$ and $c\leq c_{\rm max}$, do
4:Set $\bm{\rho}^{(c)}$, $\overline{L}^{(c)}$, and $\textbf{r}^{(c)}$ to the solution of the convex problem in (53), where the previous iterates $\bm{\rho}^{(c-1)}$, $\overline{L}^{(c-1)}$, and $\textbf{r}^{(c-1)}$ are taken as constant.
5:If $\chi_{0}<\epsilon_{\chi}$, set $\chi_{0}=0$ for the next iteration.
6:End while
7:Output: The transmit power coefficients $\bm{\rho}^{(c)}$ and the blocklength $L^{(c)}$.
VI-B Minimizing E2E Energy Consumption
In this subsection, our objective is to minimize the E2E energy consumption. The optimization problem is formulated as follows
@@ VI-B Minimizing E2E Energy Consumption
The total power consumption can be rewritten as
$P_{\rm total}=P_{\rm FIXED}+\Delta^{\rm tr}\|\bm{\rho}\|^{2}+\frac{1}{L}f_{1}+\frac{L_{d}}{L}f_{2}$ (55)
where
$P_{\rm FIXED}=\sum_{k=1}^{N_{\rm tx}}P_{\textrm{AP},0}^{\rm tx}+\sum_{r=1}^{N_{\rm rx}}P_{\textrm{AP},0}^{\rm rx}+P_{\rm fixed}+\frac{P_{\rm cloud,0}^{\rm proc}}{\sigma_{\rm cool}}$ (56)
$\quad+C_{\rm se-detector}\Bigg{)}$ (57)
$f_{2}=\frac{\Delta_{\rm cloud}^{\rm proc}}{\sigma_{\rm cool}C_{\rm max}}\frac{B}{10^{9}}\left(20MN_{\rm ue}N_{\rm tx}+12MN_{\rm tx}+C_{{\rm se-prep},m}\right).$ (58)
Similar to the approach of problem P1, we define a new variable $\bar{L}$ where $L-L_{p}\leq\frac{1}{\bar{L}}$. Then, the objective function (54a) is equivalent to
$\mathsf{F_{2}}\triangleq\frac{1}{B}\left(L\left(P_{\rm FIXED}\right)+\left(L-L_{p}\right)f_{2}+\Delta^{\rm tr}\frac{\|\bm{\rho}\|^{2}}{\bar{L}}\right).$ (59)
The non-convex optimization problem in (54) can be solved using the same approach in P1 as follows:
$\underset{\bm{\rho},\bm{\chi},\textbf{r}\geq\textbf{0},L,\bar{L}>L_{p},\chi_{0}\geq 0}{\textrm{minimize}}\quad\mathsf{F_{2}}+\lambda\chi_{0},$ (60a)

### v3

@@ V-B GOPS Analysis of Digital Operations for Sensing at the Cloud
In this subsection, we present the GOPS analysis for sensing operations, including signal transmission and processing the received signal. The GOPS corresponding to the sensing transmissions consists of computing the sensing precoding vector $\textbf{w}_{0}$, denoted by $C_{\rm prec-comp}^{\rm s}$, and obtaining the sensing signal by multiplying the sensing symbols by the sensing precoding vector and the sensing power coefficient. Multiplying the sensing symbols by the sensing precoding vector and the sensing power coefficient costs $12L_{d}MN_{\rm tx}$ real multiplications/divisions. The unitary matrix U in (6) (subspace spanned by the UE channel estimation vectors) is already obtained when computing the RZF precoding vectors by matrix inversion and the corresponding $\mathbf{L}\mathbf{D}\mathbf{L}^{H}$ decomposition [37, Lem. B.2]. Therefore, the computational complexity of the ZF precoding vector $\textbf{w}_{0}$ is given by
@@
After transmitting the signal in downlink, the reflected signals along with the interference signals are received at the receiver APs and sent to the cloud. At the cloud, these signals are processed for a specific sensing application. In this work, we consider target detection and assess the number of real multiplications/divisions required to compute the test statistics. For target detection problems, we usually compute the test statistics and compare them with a threshold. The target is declared detected if the value of the test statistics is greater than the threshold. We assume that the threshold is constant. Therefore, we can neglect the computational complexity of obtaining the threshold. However, test statistics should be obtained for each transmission. Hence, the sensing GOPS is obtained as
@@
where $C_{\rm se-comp}$ denotes the computational complexity for computing the detector test statistics.
Given the clutter-unaware detector in (11), the computational complexity is computed as
@@
where we used the block diagonal structure of C in getting the inverse of it. The computational complexity for each step of computing the test statistics of clutter-aware detector in (12) are listed in Table II.
@@ VI Joint blocklength and Power Optimization
We aim to jointly optimize the blocklength and the power control coefficients to minimize the E2E energy consumption while URLLC and sensing requirements are satisfied.  [^3] In addition to considering radio and cloud processing energy consumption, it is important to take fronthaul energy consumption into account when assessing E2E energy consumption. However, in our analysis, where the primary focus is on the interplay between processing and radio resources, we treat fronthaul energy consumption as a fixed component and do not include it in our considerations. The corresponding algorithm to this optimization problem is represented by E2E JPB-SeURLLC+. The optimization problem is formulated as follows:
$\underset{\bm{\rho}\geq\textbf{0},L>L_{p}}{\textrm{minimize}}\quad$ $E_{\rm total}=\frac{L_{d}}{B}\Delta^{\rm tr}\|\bm{\rho}\|^{2}+\frac{L}{B}\left(P_{\rm FIXED}+\frac{1}{L}f_{1}+\frac{L_{d}}{L}f_{2}\right)\,$ (40a)
subject to $\epsilon^{\rm(ub)}_{i}\leq\epsilon^{\rm(th)}_{i},\quad\forall i>0$ (40b)
$L\leq L_{\rm max,c},\quad$ (40c)
$R_{\rm s}\geq R_{\rm s}^{\rm th},\quad$ (40d)
$\overline{\mathsf{SINR}}_{\rm s}\geq\gamma_{\rm s},\quad$ (40e)
$P_{k}\leq P_{\rm tx},\quad\quad k=1,\ldots,N_{\rm tx}$ (40f)
where
$P_{\rm FIXED}=\sum_{k=1}^{N_{\rm tx}}P_{\textrm{AP},0}^{\rm tx}+\sum_{r=1}^{N_{\rm rx}}P_{\textrm{AP},0}^{\rm rx}+P_{\rm fixed}+N_{\rm GPP}\frac{P_{\rm cloud,0}^{\rm proc}}{\sigma_{\rm cool}}$ (41)
$\quad+C_{\rm se-detector}\Bigg{)}$ (42)
$f_{2}=\frac{\Delta_{\rm cloud}^{\rm proc}}{\sigma_{\rm cool}C_{\rm max}}\frac{B}{10^{9}}\left(20MN_{\rm ue}N_{\rm tx}+12MN_{\rm tx}+C_{{\rm se-prep}}\right),$ (43)
where $C_{\rm se-detector}$ is the summation of the terms in $C_{\rm se-comp}$ that does not have the multiplying factor of $L_{d}$ as it is included in $f_{1}$. On the other hand, $C_{\rm se-prep}$ includes the terms with $L_{d}$ appearing in $f_{2}$. Constraints (40b) and (40c) correspond to the URLLC requirements. Constraints (40d) and (40e) correspond to the sensing requirements, where $R_{s}^{\rm th}$ is the refreshing rate threshold and $\gamma_{\rm s}$ is the required sensing SINR that is selected according to the target detection performance requirement and $P_{\rm tx}$ is the maximum transmit power per AP. We define a new variable $\bar{L}$ where $L_{d}=L-L_{p}\leq\frac{1}{\bar{L}}$. Then, the objective function (40a) is equivalent to
$\mathsf{F}\triangleq\frac{1}{B}\left(LP_{\rm FIXED}+\left(L-L_{p}\right)f_{2}+\Delta^{\rm tr}\frac{\|\bm{\rho}\|^{2}}{\bar{L}}\right).$ (44)
@@
$\underset{\bm{\rho},\bm{\chi},\textbf{r}\geq\textbf{0},L>L_{p},\bar{L}>0}{\textrm{minimize}}\quad\mathsf{F}$ (45a)
@@
$\underset{\bm{\rho},\bm{\chi},\textbf{r}\geq\textbf{0},L>L_{p},\bar{L}>0,\chi_{0}\geq 0}{\textrm{minimize}}\quad\mathsf{F}+\lambda\chi_{0}$ (46a)
$\textrm{subject to}\quad{L-L_{p}}\leq\frac{2}{\overline{L}^{(c-1)}}-\frac{\overline{L}}{\left(\overline{L}^{(c-1)}\right)^{2}},$ (46b)
@@
The steps of the FPP-SCA and CCP procedure, E2E JPB-SeURLLC+ algorithm, are outlined in Algorithm 1. We empirically observed that setting $\overline{L}^{(0)}=\frac{1}{L_{\rm max}-L_{p}}$, $\rho_{0}^{(0)}=0$ and $\sqrt{\rho_{i}}^{(0)}=10^{-3}\sqrt{P_{\rm tx}/N_{\rm ue}}$ for $i>0$ yields satisfactory results.
Algorithm 1 FPP-SCA and CCP Procedure for Solving Problem (45), E2E JPB-SeURLLC+ algorithm
1:Initialization: Initialize $\rho_{0}^{(0)}=0$, $\sqrt{\rho_{i}}^{(0)}=10^{-3}\sqrt{P_{\rm tx}/N_{\rm ue}}$ for $i>0$, $\textbf{r}^{(0)}\geq\textbf{0}$, and blocklength $\overline{L}^{(0)}=\frac{1}{L_{\rm max}-L_{p}}$. Set solution accuracy parameters $\epsilon,\epsilon_{\chi}>0$, and $\lambda>0$. Set iteration counter $c=0$, maximum number of iterations $c_{\rm max}$, the initial objective value $\mathsf{F}^{(0)}=\infty$ and define the improvement metric as $\Delta\mathsf{F}^{(c)}=\mathsf{F}^{(c-1)}-\mathsf{F}^{(c)}$, and $\Delta\mathsf{F}^{(0)}=\infty$.
@@
3:while $\Delta\mathsf{F}^{(c-1)}\geq\epsilon$ and $c\leq c_{\rm max}$ do
4:  Solve the convexified problem in (46) using the previous iterates $\bm{\rho}^{(c-1)}$, $\overline{L}^{(c-1)}$, and $\textbf{r}^{(c-1)}$ as constants, and update $\bm{\rho}^{(c)}$, $\overline{L}^{(c)}$, and $\textbf{r}^{(c)}$ accordingly.
5:  if $\chi_{0}<\epsilon_{\chi}$ then
6:   Set $\chi_{0}=0$ for the next iteration.
7:  end if
8:  $c\leftarrow c+1$
9:end while
10:Output: Transmit power coefficients $\bm{\rho}^{(c)}$ and the optimized blocklength $L^{(c)}$.
VII Numerical Results
@@
The large-scale fading coefficients, shadowing parameters, probability of LOS, and the Rician factors are simulated based on the 3GPP Urban Microcell model, defined in [47, Table B.1.2.1-1, Table B.1.2.1-2, Table B.1.2.2.1-4]. The path losses for the Rayleigh fading target-free channels are also modeled by the 3GPP Urban Microcell model with the difference that the channel gains are multiplied by an additional scaling parameter equal to 0.30.30.3 to suppress the known parts of the target-free channels due to LOS and permanent obstacles [40]. The sensing channel gains are computed by the bi-static radar range equation [7]. The carrier frequency, the bandwidth, and the noise variance are set to 1.91.91.9 GHz, 200200200 KHz, and 114-114- 114 dBm, respectively. The number of pilot symbols is $L_{p}=10$. The regularization parameter $\delta$ in (5) is set to the noise variance.
The spatial correlation matrices for the communication channels are generated by using the local scattering model in [34, Sec. 2.5.3]. The RCS of the target is modeled by the Swerling-I model with $\sigma_{\rm rcs}=0$ dBsm. For all the UEs, the packet size, maximum transmission delay, and DEP threshold are $b_{i}=256$ bits, $D_{i}^{\rm(th)}=1$ ms, and $\epsilon_{i}^{\rm(th)}=10^{-5}$, respectively. The sensing SINR threshold is $\gamma_{s}=0$ dB, the false alarm probability threshold is $P_{\rm fa}=0.03$, and the refreshing rate threshold is $R_{\rm s}^{\rm th}=10$ updates per second [6, Table 6.2-1], unless otherwise stated. The remaining parameters are detailed in Table III, where the values are consistent with those in [18].
TABLE III: SIMULATION PARAMETERS
$\Delta^{\rm tr},\sigma_{\rm cool}$ 4, 0.9 $P_{\rm fixed}$ 120 W
$P_{\textrm{AP},0}^{\rm tx},P_{\textrm{AP},0}^{\rm rx}$ $6.8\cdot M$ W $P_{\rm cloud,0}^{\rm proc}$ 81 W
$\Delta_{\rm cloud}^{\rm proc}$ 288 W $C_{\rm max}$ 700.94 GOPS
We compare the performance of the proposed algorithm, E2E JPB-SeURLLC+, which aims to jointly optimize power and blocklength to minimize E2E energy consumption, against two benchmark schemes: (i) JPB-SeURLLC+, which minimizes only the transmission energy consumption, and (ii) E2E JPB-URLLC, which targets E2E energy minimization in URLLC systems without sensing integration. We also evaluate performance under both clutter-aware and clutter-unaware MAPRT detectors, as defined in (11) and (12), respectively. Our analysis focuses on the total energy consumed to complete one cycle of communication and sensing tasks. The communication task requires satisfying a maximum DEP threshold for each UE, while the sensing task involves detecting a target at a candidate location with a specified false alarm probability threshold. The total E2E energy includes ISAC transmission, sensing reception, and processing components for both tasks.
Figure 3: Convergence condition for E2E JPB-SeURLLC+, JPB-SeURLLC+, and E2E JPB-URLLC algorithms with clutter-aware detector and $N_{\rm rx}=2$.
Figure 4: (a) Detection probability, (b) transmission energy, and (c) E2E energy vs. sensing SINR threshold with $N_{\rm rx}=2$.
Figure 5: Maximum blocklength threshold and network availability vs. refreshing rate threshold for $N_{\rm rx}=2$ and the E2E JPB-SeURLLC+ algorithm.
Figure 6: (a) Detection probability, (b) transmission, and (c) total energy consumption vs. DEP threshold for $N_{\rm rx}=2$, $\gamma_{\rm s}=0$ dB, and $R_{\rm s}^{\rm th}=1000$ updates/s.
Figure 7: Breakdown of energy consumption for completing one sensing and communication task with clutter-aware detector, $N_{\rm rx}=2$, and $\gamma_{\rm s}=0$ dB.
Figure 8: (a) Computational complexity for one sensing task, (b) sensing GOPS, (c) number of required GPPs vs. blocklength.
(a)(b)
Figure 9: (a) Number of antenna elements per AP, and (b) total energy consumption for a minimum detection probability of 0.90.90.9 with E2E JPB-SeURLLC+ algorithm and $N_{\rm rx}=1,2$.
Fig. 3 illustrates the convergence behavior of the proposed and benchmark algorithms. The objective of the E2E algorithms is to minimize the total energy consumption, whereas the JPB-SeURLLC+ algorithm focuses solely on minimizing transmission energy consumption. In the figure, the values for JPB-SeURLLC+ are scaled by $10^{5}$ since they are significantly lower compared to the total energy values. In the algorithms, the solution accuracy parameters are set as $\epsilon=10^{-3}$, $\epsilon_{\chi}=10^{-6}$, $\lambda=10$, and the maximum iteration is set to $c_{\rm max}=30$.
In the following, we first compare the performance of the proposed E2E JPB-SeURLLC+ algorithm with relevant benchmarks, followed by a comprehensive analysis of the proposed algorithm under various communication and sensing requirements. Then, we investigate the interplay between the number of receive APs, the number of antenna elements per AP, and the complexity level of the detector to balance the sensing performance and energy consumption in the system. Detection probability results are obtained empirically through Monte Carlo simulation.
VII-A Impact of Sensing and Communication Requirements
The effect of the sensing SINR threshold $\gamma_{\rm s}$ is evaluated in Fig. 4 for $N_{\rm rx}=2$. As shown in Fig. 4(a), detection probability generally increases with $\gamma_{s}$, but performance varies based on the detector and resource allocation strategy. Clutter-aware detectors consistently outperform clutter-unaware ones due to their advanced signal processing. As this detector estimates the target-free channels, its performance does not change significantly with increasing $\gamma_{\rm s}$ while the performance of clutter-unaware detector accordingly. The proposed E2E JPB-SeURLLC+ with a clutter-aware detector achieves the highest detection probability up to 0.980.980.98, while the same algorithm with a clutter-unaware detector reaches only detection probability of 0.630.630.63. The JPB-SeURLLC+ algorithm that prioritizes minimizing transmission energy consumption, as opposed to E2E energy minimization, tend to result in lower detection performance. This is because the E2E algorithms try to minimize the total energy consumption by reducing the blocklength. This approach allows higher transmission power levels–as shown in Fig. 4(b)– to meet the requirements, which in turn enhance detection capability. In contrast, E2E JPB-URLLC fails to exceed 0.5 detection probability, even with a clutter-aware detector—highlighting the need to jointly optimize sensing and communication in URLLC systems.